Write the equation of a line given a slope of one fourth and a point (2, -4)
Answer by Alan3354(69443)

Write the equation of a line given a slope of one fourth and a point (2, -4)
-------------
Use y-y1 = m*(x-x1) where m = slope and (x1,y1) is the point.
y+4 = (1/4)*(x-2)